# Active Search for Low-altitude UAV Sensing and Communication for Users at Unknown Locations

### v1

@@ V-A Environment Setup and Scenarios
Two application scenarios are evaluated in our experiments. For a sum-rate application, one UAV is placed to establish LOS relay channels for $K$ ground users and a BS under decode-and-forward relaying. Consider the path loss model of millimeter wave cellular reported in [28, 29], where the deterministic LOS channel gain $g_{k}(\mathbf{x})$ is modeled as $g_{k}(\mathbf{x})=46.53+20.0\log_{10}d(\mathbf{x},\mathbf{u}_{k})$. The variance $\sigma^{2}$ of measurement uncertainty $\xi$ in (4) is set as $5$ dB. It is assumed that the mmW beam alignment has been done for every UAV position. For a balancing application, one UAV is placed to provide detection services for eight sensing targets and maintains a backhaul communication link with a BS simultaneously. Particularly, the objective functions for sensing are specified as SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=p_{k}g_{k}(\mathbf{x})/(\bar{N}_{0}W)$ where the noise power density $\bar{N}_{0}$ is set as $-150$ dBm/Hz, and the bandwidth $W$ is set as $1$ GHz.
@@ V-B UAV-assisted Communication and Sensing
Figure 7: System capacity versus transmit power of UAV
@@
Figure 8: System capacity versus number of users
Figure 9: CDF versus $\text{SNR}=\min_{k\in\mathcal{K}}\{p_{k}^{*}g_{k}(\mathbf{x})/N_{0}\}$ in a UAV-assisted sensing and communication application.
@@ VI Conclusion
In this paper, we developed an efficient online trajectory for UAV placement without prior knowledge of user locations, channel model parameters, or terrain structure. We analytically characterized the equipotential surface and proposed an LOS discovery strategy on it, utilizing perturbation theory to guide the UAV search direction. Furthermore, we introduced spiral trajectories to construct local channel models in the LOS regime using local polynomial regression, independent of user locations. After deriving the optimal measurement pattern, we minimized the MSE of the locally estimated channel gain and determined the optimal measurement range. Experimental results on real urban maps demonstrated that our approach achieves over $95\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search, offering a practical solution for UAV-assisted sensing and communication in complex environments.
@@ Appendix D Proof of Theorem 2
Next, the bounds of $\mathbb{V}\{\hat{g}(\mathbf{x})\}$ and $|\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}|$ with $r_{0}=d(\mathbf{x},\mathbf{c}_{0})$ are derived as below.
@@
In addition, as $d(\mathbf{x}_{m},\mathbf{c}_{0})\leq r_{1}$ and $d(\mathbf{x},\mathbf{c}_{0})\leq r_{0}$, the Lipschitz condition implies that $|e(\mathbf{x}_{m},\mathbf{c}_{0})|\leq\frac{L_{g}^{2}}{2}d(\mathbf{x}_{m},\mathbf{c}_{0})^{2}\leq L_{g}r_{1}^{2}/2$ and $|e(\mathbf{x},\mathbf{c}_{0})|\leq L_{g}r_{0}^{2}/2$. The bias can be upper bounded as
@@
Appendix E Supplementary Materials 1: Complexity Analysis
As this paper studies trajectory design for UAV online search, the following two metrics are essential for understanding the complexity: the trajectory length and the per-step computational complexity every time the UAV adjusts its course.
Proposition 3 derives the upper bound of the search trajectory under certain conditions in a balancing problem.
Proposition 3 (Upper bound of trajectory length).
Suppose that the channel gain satisfies $g_{k}(\mathbf{x})=b_{0}-10\log_{10}(d^{2}(\mathbf{x},\mathbf{u}_{k}))$. For the balancing problem defined in (6) with $f_{k}(g_{k}(\mathbf{x}),p_{k})=\log_{2}(1+p_{k}g_{k}(\mathbf{x}))$, the trajectory length $L$ is found to be linear in some geometry parameters of the environment:
$L\leq\pi(H_{0}+R)\sqrt{3\pi^{2}r_{1}^{2}+1}$ (52)
where $H_{0}$ is the initial search altitude, $R$ is the radius of the equipotential surface given in (11), and $r_{1}$ is the measurement radius.
Proof.
@@ Appendix E Supplementary Materials 1: Complexity Analysis
Phase 2: trajectory in the non-full-LOS region. The trajectory in the non-full-LOS region satisfies $f_{0}(g_{0}(\mathbf{x}_{\text{s}}(t)))=C$. Therefore, ${\bf x}_{\text{s}}(t)$ in the non-full-LOS region is an arc with a radius smaller than or equal to $H_{0}$ and a central angle less than or equal to $\pi$. Then, the trajectory $\mathbf{x}_{\text{s}}(t)$ in the non-full-LOS region is upper bound by $\pi H_{0}$.
@@
Finally, we combine the LOS discovery trajectory ${\bf x}_{\text{s}}(t)$ and the spiral measurement trajectory ${\bf x}_{\text{r}}(t)$. Given the upper bound of ${\bf x}_{\text{s}}(t)$ as $\pi(H_{0}+R)$, the total search time is upper bound by $T\leq\pi(H_{0}+R)/v$. According to the design of the spiral trajectory in Section III-E, let $\omega=4\pi/M$ and $v=2/\sqrt{M^{2}-1}$. Then, the total length $L$ of the superposed trajectory is upper bound by
@@
Proposition 3 implies that reducing the initial search altitude is likely to reduce the search trajectory length $L$ as it results in a smaller upper bound of $L$. In addition, the radius $R$ of the equipotential surface given in (11) consists of the distance between the user cluster and the BS. Consider a special case where ${\bf u}_{0}=[0,0,0]^{\text{T}}$, $K=1$, $P_{0}\neq P_{\text{T}}$, and $P_{0}>P_{\text{T}}>0$. Then, $R^{2}$ in (11) can be simplified as
$R^{2}=\frac{P_{0}}{P_{0}-P_{\text{T}}}\left(\frac{P_{0}}{P_{0}-P_{\text{T}}}-1\right)d^{2}({\bf u}_{0},{\bf u}_{1}).$ (56)
The equation (56) indicates that the radius $R$ is monotonically increasing w.r.t. the distance $d({\bf u}_{0},{\bf u}_{1})$ between the user and the BS. Thus, the trajectory length is also related to $d({\bf u}_{0},{\bf u}_{1})$. Furthermore, the power allocation also affects the shape of the equipotential surface. For example, when the UAV transmit power $P_{\text{T}}$ is much smaller than the BS transmit power $P_{0}$, the radius $R$ of the equipotential surface in (56) will be quite small.
As the proposed scheme is an online strategy implemented on UAV s, the per-step computational complexity is also derived. Specifically, the computational complexity of the proposed algorithm mainly comes from the local channel construction in Step 3), and the update of UAV positions in Step 4) in Algorithm 1.
  • Computational complexity of Step 3): The local channel construction in Step 3) builds a local channel model for each user using the closed-form formula in equation (17), and thus, the computational complexity is given by $O(KM)$, where recall that $K$ is the number of users and $M$ is the number of measurements used for local channel model construction.
  • Computational complexity of Step 4): The update of UAV positions in Step 4) has a computational complexity of $O((K+N)^{3})$, where $N$ is the number of constraints for resource allocation. Step 4) depends on closed-form formulations in (14), (37), (38), (39), (40), (42), (43) and (44), and the major complexity comes from the matrix inverse operation in (37).
Appendix F Supplementary Materials 2: Proof of Lemma 1
@@
Lemma 2 (Eigenvalue inequality [32, Corollary 4.3.12]).
@@
Lemma 3 (Trace inequality [32, Equation 2.4.2.1]).

### v4

@@ V-A Environment Setup and Scenarios
Two application scenarios are evaluated in our experiments. For a sum-rate application, one UAV is placed to establish LOS relay channels for $K$ ground users and a BS under decode-and-forward relaying. Consider the path loss model of millimeter wave cellular reported in [29, 30], where the deterministic LOS channel gain $g_{k}(\mathbf{x})$ is modeled as $g_{k}(\mathbf{x})=46.53+20.0\log_{10}d(\mathbf{x},\mathbf{u}_{k})$. The variance $\sigma^{2}$ of measurement uncertainty $\xi$ in (4) is set as $5$ dB. It is assumed that the mmW beam alignment has been done for every UAV position. For a balancing application, one UAV is placed to provide location estimation services for $8$ sensing targets and maintain a backhaul communication link with a BS simultaneously. Particularly, the objective functions for sensing are specified as SNR, i.e., $f_{k}(g_{k}(\mathbf{x}),p_{k})=p_{k}g_{k}(\mathbf{x})/(\bar{N}_{0}W)$ where the noise power spectral density $\bar{N}_{0}$ is set as $-150$ dBm/Hz, and the bandwidth $W$ is set as $1$ GHz.
@@ V-B UAV-assisted Communication and Sensing
Figure 7: System capacity versus transmit power of UAV in a sum-rate application
@@
Figure 8: System capacity versus number of users in a sum-rate application
Figure 9: CDF versus SNR of the weakest sensing link in a balancing application.
Fig. 10 shows two examples of convergence process of the proposed scheme. The velocity of the UAV is set as $10$ meters per second. It is observed that the system capacity increases with the search time because the proposed scheme searches in the direction of increasing the objective value in the full-LOS region or maintaining the current objective value in the non-full-LOS region.
Define the convergence time as the search time at which the system performance no longer improves, in which case, we consider that the best UAV position is discovered. Fig. 11 summarizes the convergence time versus number of users. It is observed that the convergence time decreases with the increase of number of users $K$. The reason is that the radius of the equipotential surface decreases with the increase of $K$, which is analytically shown in Proposition 2. Since the maximum height of the equipotential surface is smaller, the trajectory length will be shorter as it takes less time to reach the minimum flight height.
Figure 10: Convergence process: system capacity versus search time.
Figure 11: The search time to discover the best UAV position versus the number of users. The shaded area represents a possible search time realization during our $2000$ simulations.
@@ VI Conclusion
This paper developed an efficient online trajectory for optimal UAV placement without prior knowledge of user locations, channel model parameters, and terrain structure. We analytically characterized the equipotential surface and proposed an LOS discovery trajectory on it, utilizing perturbation theory to guide the UAV search direction. Additionally, we developed a class of spiral trajectories to construct a local channel map in the LOS regime using local polynomial regression without depending on user locations. After deriving the optimal measurement pattern, we minimized the MSE of the locally estimated channel gain and determined the optimal measurement range. Experimental results on real urban maps demonstrated that our approach achieves over $94\%$ of the performance of a 3D exhaustive search scheme with just a $3$-kilometer search in a complex environment.
@@ Appendix D Proof of Theorem 2
Next, the bounds of $\mathbb{V}\{\hat{g}(\mathbf{x})\}$ and $|\mathbb{E}\{\hat{g}(\mathbf{x})-g(\mathbf{x})\}|$ with $r_{0}=d(\mathbf{x},\mathbf{c}_{0})$ are derived as follows.
@@
In addition, as $d(\mathbf{x}_{m},\mathbf{c}_{0})\leq r_{1}$ and $d(\mathbf{x},\mathbf{c}_{0})\leq r_{0}$, the Lipschitz condition implies that $|e(\mathbf{x}_{m},\mathbf{c}_{0})|\leq\frac{L_{g}^{2}}{2}d(\mathbf{x}_{m},\mathbf{c}_{0})^{2}\leq L_{g}r_{1}^{2}/2$ and $|e(\mathbf{x},\mathbf{c}_{0})|\leq L_{g}r_{0}^{2}/2$. The bias is upper bounded as
@@
Supplementary Materials
Appendix E Proof of Proposition 3
@@
Phase 2: trajectory in the non-full-LOS region. The trajectory in the non-full-LOS region satisfies $f_{0}(g_{0}(\mathbf{x}_{\text{s}}(t)))=C$. Therefore, ${\bf x}_{\text{s}}(t)$ in the non-full-LOS region is an arc with a radius smaller than or equal to $H_{0}$ and a central angle less than or equal to $\pi$. Then, the trajectory $\mathbf{x}_{\text{s}}(t)$ in the non-full-LOS region is upper bounded by $\pi H_{0}$.
@@
Finally, we combine the LOS discovery trajectory ${\bf x}_{\text{s}}(t)$ and the spiral measurement trajectory ${\bf x}_{\text{r}}(t)$. Given the upper bound of ${\bf x}_{\text{s}}(t)$ as $\pi(H_{0}+R)$, the total search time is upper bounded by $T\leq\pi(H_{0}+R)/v$. According to the design of the spiral trajectory in Section III-E, let $\omega=4\pi/M$ and $v=2/\sqrt{M^{2}-1}$. Then, the total length $L$ of the superposed trajectory is upper bounded by
@@
Appendix F Per-step Computational Complexity Analysis
The computational complexity of the proposed algorithm mainly comes from the local channel construction in Step 3, and the update of UAV positions in Step 4 in Algorithm 1.
  • Computational complexity of Step 3: The local channel construction in Step 3 builds a local channel model for each user using the closed-form formula in equation (17), and thus, the computational complexity is given by $O(KM)$, where recall that $K$ is the number of users and $M$ is the number of measurements used for local channel model construction.
  • Computational complexity of Step 4: The update of UAV positions in Step 4 has a computational complexity of $O((K+N)^{3})$, where $N$ is the number of constraints for resource allocation. Step 4 depends on closed-form formulations in (14), (37), (38), (39), (40), (42), (43) and (44), and the major complexity comes from the matrix inverse operation in (37).
Appendix G Proof of Lemma 1
@@
Lemma 2 (Eigenvalue inequality [33, Corollary 4.3.12]).
@@
Lemma 3 (Trace inequality [33, Equation 2.4.2.1]).